OpenStudy (anonymous):

Given that U is the centroid of ΔOPQ, find PS. A.1 B. 1.8 C. 3.6 D. 5.4

OpenStudy (anonymous):

 PLS PLS PLS help me understand this!!! HELP! :(

OpenStudy (anonymous):

@uri do you know geometry??

ganeshie8 (ganeshie8):

3.6 = 2(x+.8)

ganeshie8 (ganeshie8):

u still on track ?

OpenStudy (anonymous):

im still here. o:

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

!

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

PS = PU + US = 3.6 + 1+0.8 = 3.6 + 1.8

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

where did you get x=1 ? 5.4

ganeshie8 (ganeshie8):

u oly said, u got x=1 ?

OpenStudy (anonymous):

5.4 is the answer.

OpenStudy (anonymous):

and i asked how you got x=1 ? where did tht come from? o:

ganeshie8 (ganeshie8):

go thru ur replies above, u told me, u solved x = 1

OpenStudy (anonymous):

oh yeah! ok

ganeshie8 (ganeshie8):

u shud knw how u got x = 1 i just trusted wat u said, and went ahead :|

OpenStudy (anonymous):

its been a long day... sorry

ganeshie8 (ganeshie8):

its okay

OpenStudy (anonymous):

lol thx, think you could help with a simpler problem?

ganeshie8 (ganeshie8):

shoot